Ms. Jordan has been given a loan of $2500 for 1 year. If
the interest charged is $275, what is the interest rate on the loan?

Answer by priya(51)

let P be the principle ,R be the interest rate (%),N be the number of years,
then the formula is given by,
I = P*N*R/100
= 2500 * 1* R/100
HERE I IS GIVEN ,THAT IS 275,

275 = 25 *1* R (Simplyfying)
Therefore R = 275 /25
= 11 %
this is the answer.
VISIT MY SITE TO LEARN MATHS ONLINE!!!
PRIYA